# From Forecast to Action: Uncertainty-Aware UAV Deployment for Ocean Drifter Recovery

### v1

@@ 1. INTRODUCTION
Maritime accidents can cause severe damage to both human life and the environment, making rapid response crucial. According to maritime accident statistics, the incidence of such accidents in South Korea remains high (choe2025development). In response, some studies have focused on predicting the trajectories of ocean drifters (kim2024evolutionary; jurkus2023application; li2023ship). Other studies have addressed the deployment optimization of unmanned aerial vehicles (UAVs) to detect drifting particles (lee2024memetic; hong2024memetic). Although recent advances in machine learning–based trajectory prediction have achieved practical levels of accuracy, these predictions have rarely been effectively integrated into UAV deployment strategies. To the best of our knowledge, no prior work has proposed an end-to-end framework that simultaneously predicts drifting trajectories and optimizes UAV deployment decisions. Although trajectory prediction has been widely studied, its integration into downstream deployment strategies remains limited. Similarly, UAV deployment research has extensively used meta-heuristic algorithms, yet the specific design of cost functions is often insufficiently considered (elmachtoub2022smart).
@@
While predict-then-optimize frameworks are well established for integrating forecasting with decision-making (kotary2023predict; sadana2025survey; vanderschueren2022predict), most existing studies have focused on relatively small-scale problems, such as the shortest path or portfolio optimization, where exact optimization methods like linear programming (wilder2019melding), quadratic programming, or mixed-integer programming (MIP) (elmachtoub2022smart) are applicable. However, applying these frameworks to large-scale, dynamic problems such as UAV deployment for maritime search and rescue poses unique challenges. The vast and continuous search space involved in optimizing the placement of multiple UAVs under uncertainty quickly renders exact methods computationally intractable, making them impractical for time-critical search and rescue (SAR) operations. Consequently, meta-heuristic algorithms are required to efficiently explore the enormous solution space and obtain high-quality solutions.
@@
In this study, we propose a novel predict-then-optimize framework that integrates ocean drifter trajectory prediction with UAV deployment optimization, enabling robust and intelligent maritime rescue operations. We employ a language model to forecast drifter trajectories and use simulated annealing (SA), particle swarm optimization (PSO), and genetic algorithm (GA) to optimize UAV deployment around Gaussian-based particle samples near the predicted paths, accounting for potential prediction errors. To evaluate UAV deployment effectiveness, we introduce a custom metric. While prior studies have demonstrated the reliability of language model-based trajectory prediction (kim2024evolutionary), our work emphasizes the integration of these predictions with an optimized UAV deployment strategy. The framework is validated using real-world oceanographic data. Our main contributions can be summarized as follows:
@@ 2.2.2. Dynamic Detection Radius
The UAV model used in this study is the DJI Matrice 300 RTK. While its base detection radius is 100 meters, we model an effective search coverage that extends up to 600 meters, assuming the UAV follows a spiral search pattern with up to three revolutions, as illustrated in Figure 2. This operational model is constrained by the UAV’s maximum flight time of 55 minutes and travel speed of 1.38 km/min, within which it must reach its assigned coordinate, perform the search, and return (hao2024monitoring).
@@ 2.2.3. Probability of Detection (PoD) Model
While the geometric coverage is used as the fitness function during optimization for computational efficiency, the final evaluation of a deployment strategy considers the probability of detection. Since the UAV performs a spiral motion in a specific direction, and the ocean drifter may approach from the opposite direction, detection is not guaranteed. Following established SAR literature (guard2002us; agbissoh2019decision), the probability of detection (PoD) is modeled as the ratio of the dynamic detection radii to the maximum possible detection radii, expressed as:
@@ 2.3. Prediction
We used language model to predict the trajectory of an ocean drifter. While previous studies (haldar2025wide; bayindir2024lagrangian; kim2024evolutionary) have demonstrated the effectiveness of various machine learning models, recent advancements in sequence modeling have highlighted the superior performance of transformer-based architectures in capturing long-term dependencies in time series data (su2025systematic; kang2024transformer; xu2024survey).
For the prediction part of our predict-then-optimize framework, we therefore leverage a large language model (LLM) to estimate the future position of the ocean drifter, $\hat{y}_{t}$, at time step $t$. Specifically, we employ Chronos, a pretrained foundation model that reframes time series forecasting as a language modeling problem (ansari2024chronos). This approach allows the model to leverage knowledge learned from a vast and diverse corpus of time series data for robust zero-shot prediction (ansari2024chronos; wang2024deep). This zero-shot prediction capability is particularly crucial in real-world rescue scenarios, where the time-critical environment does not allow for online model training.
@@
Our prediction strategy follows an iterative forecasting strategy, which is known as a one-step-ahead recursive approach with an expanding window (dong2024temporal). At each time step, the model predicts only the next value, and the prediction is then fed back as part of the input for subsequent forecasts. Specifically, for the first prediction at time step $t=1$, the model uses the initial coordinate $y_{0}$ to forecast $\hat{y}_{1}$. To predict the coordinate at $t=2$, the model is provided with the updated historical sequence $\{y_{0},y_{1}\}$. This process is repeated, so that the prediction at any future time step $t_{p}$ relies on all available past observations $\{y_{0},y_{1},\dots,y_{p-1}\}$.
However, such an iterative forecasting strategy may accumulate errors over time. Even seemingly minor deviations can result in large spatial discrepancies in practice. Notably, in the real world, even a seemingly small difference—such as 0.1 degrees in latitude or longitude, which might appear insignificant in evaluation metrics such as mean squared error (MSE) can translate into several kilometers of physical distance. To address this, we introduced Gaussian-based particle sampling centered around the LLM predictions to better reflect spatial uncertainty and the real-world implications of prediction errors. Spatial uncertainty was modeled by sampling $k$ particles from a bivariate Gaussian distribution centered at the predicted position $\hat{y}_{t}$. In (brettin2025uncertainty), the prediction uncertainty in oceanographic applications is often assumed to follow a Gaussian distribution. The covariance matrix was defined as $\Sigma=\sigma^{2}I$, where $\sigma$ was set to four times the Haversine distance between the predicted and actual positions at time $t_{p-1}$. Each particle $p_{k}$ represents a possible location of the ocean drifter, sampled as $p_{k}\sim\mathcal{N}(\hat{y}_{t},\Sigma)$.
@@ 2.4.2. Genetic Operator
From the $N$, $N/2$ pairs are selected randomly, and genetic operator such as crossover and mutation were applied to generate $N$ offspring (aldana2024moody). We used the blend crossover (BLX-$\alpha$) operator (ranjan2025threshold), which generates offspring within an extended interval around the parent genes. Offspring $o=\{o_{1},o_{2},\ldots,o_{n}\}$ is generated from parents $x=\{x_{1},x_{2},\ldots,x_{n}\}$ and $y=\{y_{1},y_{2},\ldots,y_{n}\}$, where each $o_{i}$ is uniformly randomly chosen from the interval $\min(x_{i},y_{i})-\alpha D,\ \max(x_{i},y_{i})+\alpha D]$, with $D=|x_{i}-y_{i}|$ and $\alpha$ is constant.
@@ 2.4.3. Particle Swarm Optimization
We also implemented particle swarm optimization (PSO), another population-based meta-heuristic inspired by the social behavior of bird flocking or fish schooling (bonyadi2017particle). A swarm of $N$ particles is initialized, where each particle represents a complete UAV deployment solution $u=\{(x_{1},y_{1}),\allowbreak\ldots,\allowbreak(x_{u},y_{u})\}$. The initial position of each UAV for every particle is generated randomly within the search area $A_{r}$, identical to the initialization process described for the GA. Furthermore, each particle is assigned an initial velocity vector, typically set to zero or a small random value (guo2025parameters).
In each iteration, the velocity and position of each particle are updated based on its personal best known position ($p_{\text{best}}$) and the global best known position in the swarm ($g_{\text{best}}$) (guo2025parameters). The velocity $v_{i}$ and position $x_{i}$ of the $i$-th particle are updated as follows:
@@ 2.4.4. Simulated Annealing
Simulated Annealing (SA) is a global optimization algorithm inspired by the physical annealing process, a technique that involves heating and controlled cooling of a material to alter its physical properties (kirkpatrick1983optimization). SA operates by exploring new solutions in the neighborhood of the current solution (khurshid2024hybrid). When a better solution is found, the algorithm moves toward it to improve performance. This process is repeated over multiple iterations, where at iteration $k$, the cooling parameter controls the acceptance of new solutions. In the early stages, a high cooling parameter allows the algorithm to accept a wider range of solutions and explore the search space broadly. As the iterations progress, the cooling parameter gradually decreases, enabling a more refined search and guiding the algorithm toward the final solution.
@@ 48    $d_{c}\leftarrow\text{HaversineDistance}(P_{repaired}[p],C_{center})$
We employed linear interpolation, which estimates intermediate values when the endpoints are known, to relocate UAVs that are outside the $A_{r}$ area back within the allowed region (kholiavchenko2024kabr; scott2022soccertrack). Specifically, given a UAV position $p_{repaired}$, the UAV control center $C_{\text{center}}$, and the allowed search area radius $R_{\text{search}}$, the repaired UAV position $p_{\text{repaired}}$ is computed as $p_{\text{repaired}}\leftarrow C_{\text{center}}+\frac{p_{repaired}-C_{\text{center}}}{\|p_{repaired}-C_{\text{center}}\|}\times R_{\text{search}}$. This formula computes the unit vector pointing from the center to the UAV and scales it by the search radius, ensuring that the UAV is placed exactly on the boundary. The same mechanism is used both for the initial correction and as a boundary clamping step after applying repulsive forces.
We then applied the virtual force algorithm (VFA) to minimize overlap between UAVs. In VFA, virtual forces such as attractive and repulsive forces are typically employed to control the movement of agents by drawing them toward targets and pushing them away from obstacles or other agents (yoon2013efficient). However, in our study, only the repulsive force $\vec{F}_{r}$ is used to prevent UAVs from overlapping. When the haversine distance between two UAVs, $u_{i}=(x_{i},y_{i})$ and $u_{j}=(x_{j},y_{j})$, is less than the sum of their detection radii, the positions of $u_{i}$ and $u_{j}$ are updated by adding $\alpha_{r}\cdot\frac{\vec{f}_{r}}{n_{r}}$, where $\vec{f}_{r}\leftarrow\vec{f}_{r}+\left(\frac{r_{u_{i}}+r_{u_{j}}}{d(u_{i},u_{j})}\right)\cdot(u_{j}-u_{i})$, $d(u_{i},u_{j})$ is the haversine distance between $u_{i}$ and $u_{j}$, $r_{u_{i}}$ and $r_{u_{j}}$ are their detection radii, $n_{r}$ is the number of repulsive interactions, and $\alpha_{r}$ is a constant weighting parameter.
@@ 3.2. Experimental setup
Table 1 lists the settings of parameters used in this study. The population size and the number of generations in the GA were both set to 50. The crossover and mutation rates were set to 100% and 10%, respectively. We used three instances to evaluate the proposed framework. For BLX-$\alpha$, the constant parameter $\alpha$ is set to 0.5, which has been widely used (yoon2021maximizing; tsutsui2001search). In the repair phase of VFA, $\alpha_{r}$ is set to 0.9. Additionally, different settings—such as the number of particles and UAVs—were applied to assess the generalizability of the framework’s performance. During the GA process, a unit length of 100 (equivalent to 100 meters) was used to segment the straight lines between particles and the accident location. In other words, we employed a coarse approximation during the GA to optimize performance efficiently, followed by a detailed evaluation for accuracy After the GA phase, a finer unit length of 1 was used to compare the actual trajectories of the ocean drifters with the UAV deployment results, and the number of initial missing ocean drifters, $K_{0}$, was set to 100. For PSO, the inertia weight $w$ was set to 0.7, and both cognitive and social coefficients, $c_{1}$ and $c_{2}$, were set to 2.0. The SA algorithm was configured with an initial temperature of $T=1.0$, a cooling rate of $\alpha=0.95$, and a number of iterations was set to 10.
@@ 3.3.1. Prediction
To evaluate the predictive performance of the language model, we compared it with models previously employed for ocean drifter forecasting (kim2025evolutionary; kim2024evolutionary). To mimic real-world scenarios, we calculated the Haversine distance between the predicted and actual values. As shown in Table 2, while the conventional benchmark models outperformed the larger Chronos-base and Chronos-large variants, which had errors over 1.4 km, the compact Chronos-tiny model achieved superior performance, yielding a lower Haversine error (0.979 km) compared to the best-performing benchmark, Linear Regression (1.142 km).
@@ 6. Acknowledgments
This research was supported by Korea Institute of Marine Science & Technology Promotion (KIMST) funded by the Ministry of Oceans and Fisheries, Korea (RS-2022-KS221629).

### v2

@@ 1. INTRODUCTION
Maritime accidents can cause severe damage to both human life and the environment, making rapid response crucial. According to maritime accident statistics, the incidence of such accidents in South Korea remains high (Choe et al., 2025). In response, some studies have focused on predicting the trajectories of ocean drifters (Kim et al., 2024; Jurkus et al., 2023; Li et al., 2023). Other studies have addressed the deployment optimization of unmanned aerial vehicles (UAVs) to detect drifting particles (Lee and Kim, 2024; Hong and Kim, 2024). Although recent advances in machine learning–based trajectory prediction have achieved practical levels of accuracy, these predictions have rarely been effectively integrated into UAV deployment strategies. To the best of our knowledge, no prior work has proposed an end-to-end framework that simultaneously predicts drifting trajectories and optimizes UAV deployment decisions. Although trajectory prediction has been widely studied, its integration into downstream deployment strategies remains limited. Similarly, UAV deployment research has extensively used meta-heuristic algorithms, yet the specific design of cost functions is often insufficiently considered (Elmachtoub and Grigas, 2022).
@@
While predict-then-optimize frameworks are well established for integrating forecasting with decision-making (Kotary et al., 2023; Sadana et al., 2025; Vanderschueren et al., 2022), most existing studies have focused on relatively small-scale problems, such as the shortest path or portfolio optimization, where exact optimization methods like linear programming (Wilder et al., 2019), quadratic programming, or mixed-integer programming (MIP) (Elmachtoub and Grigas, 2022) are applicable. However, applying these frameworks to large-scale, dynamic problems such as UAV deployment for maritime search and rescue poses unique challenges. The vast and continuous search space involved in optimizing the placement of multiple UAVs under uncertainty quickly renders exact methods computationally intractable, making them impractical for time-critical search and rescue (SAR) operations. Consequently, meta-heuristic algorithms are required to efficiently explore the enormous solution space and obtain high-quality solutions.
@@
In this study, we propose a novel predict-then-optimize framework that integrates ocean drifter trajectory prediction with UAV deployment optimization, enabling robust and intelligent maritime rescue operations. We employ a language model to forecast drifter trajectories and use simulated annealing (SA), particle swarm optimization (PSO), and genetic algorithm (GA) to optimize UAV deployment around Gaussian-based particle samples near the predicted paths, accounting for potential prediction errors. To evaluate UAV deployment effectiveness, we introduce a custom metric. While prior studies have demonstrated the reliability of language model-based trajectory prediction (Kim et al., 2024), our work emphasizes the integration of these predictions with an optimized UAV deployment strategy. The framework is validated using real-world oceanographic data. Our main contributions can be summarized as follows:
@@ 2.2.2. Dynamic Detection Radius
The UAV model used in this study is the DJI Matrice 300 RTK. While its base detection radius is 100 meters, we model an effective search coverage that extends up to 600 meters, assuming the UAV follows a spiral search pattern with up to three revolutions, as illustrated in Figure 2. This operational model is constrained by the UAV’s maximum flight time of 55 minutes and travel speed of 1.38 km/min, within which it must reach its assigned coordinate, perform the search, and return (Hao et al., 2024).
@@ 2.2.3. Probability of Detection (PoD) Model
While the geometric coverage is used as the fitness function during optimization for computational efficiency, the final evaluation of a deployment strategy considers the probability of detection. Since the UAV performs a spiral motion in a specific direction, and the ocean drifter may approach from the opposite direction, detection is not guaranteed. Following established SAR literature (Guard, 2002; Agbissoh Otote et al., 2019), the probability of detection (PoD) is modeled as the ratio of the dynamic detection radii to the maximum possible detection radii, expressed as:
@@ 2.3. Prediction
We used language model to predict the trajectory of an ocean drifter. While previous studies (Haldar et al., 2025; Bayindir et al., 2024; Kim et al., 2024) have demonstrated the effectiveness of various machine learning models, recent advancements in sequence modeling have highlighted the superior performance of transformer-based architectures in capturing long-term dependencies in time series data (Su et al., 2025; Kang and Kang, 2024; Xu et al., 2024).
For the prediction part of our predict-then-optimize framework, we therefore leverage a large language model (LLM) to estimate the future position of the ocean drifter, $\hat{y}_{t}$, at time step $t$. Specifically, we employ Chronos, a pretrained foundation model that reframes time series forecasting as a language modeling problem (Ansari et al., 2024). This approach allows the model to leverage knowledge learned from a vast and diverse corpus of time series data for robust zero-shot prediction (Ansari et al., 2024; Wang et al., 2024). This zero-shot prediction capability is particularly crucial in real-world rescue scenarios, where the time-critical environment does not allow for online model training.
@@
Our prediction strategy follows an iterative forecasting strategy, which is known as a one-step-ahead recursive approach with an expanding window (Dong et al., 2024). At each time step, the model predicts only the next value, and the prediction is then fed back as part of the input for subsequent forecasts. Specifically, for the first prediction at time step $t=1$, the model uses the initial coordinate $y_{0}$ to forecast $\hat{y}_{1}$. To predict the coordinate at $t=2$, the model is provided with the updated historical sequence $\{y_{0},y_{1}\}$. This process is repeated, so that the prediction at any future time step $t_{p}$ relies on all available past observations $\{y_{0},y_{1},\dots,y_{p-1}\}$.
However, such an iterative forecasting strategy may accumulate errors over time. Even seemingly minor deviations can result in large spatial discrepancies in practice. Notably, in the real world, even a seemingly small difference—such as 0.1 degrees in latitude or longitude, which might appear insignificant in evaluation metrics such as mean squared error (MSE) can translate into several kilometers of physical distance. To address this, we introduced Gaussian-based particle sampling centered around the LLM predictions to better reflect spatial uncertainty and the real-world implications of prediction errors. Spatial uncertainty was modeled by sampling $k$ particles from a bivariate Gaussian distribution centered at the predicted position $\hat{y}_{t}$. In (Brettin et al., 2025), the prediction uncertainty in oceanographic applications is often assumed to follow a Gaussian distribution. The covariance matrix was defined as $\Sigma=\sigma^{2}I$, where $\sigma$ was set to four times the Haversine distance between the predicted and actual positions at time $t_{p-1}$. Each particle $p_{k}$ represents a possible location of the ocean drifter, sampled as $p_{k}\sim\mathcal{N}(\hat{y}_{t},\Sigma)$.
@@ 2.4.2. Genetic Operator
From the $N$, $N/2$ pairs are selected randomly, and genetic operator such as crossover and mutation were applied to generate $N$ offspring (Aldana-Martín et al., 2024). We used the blend crossover (BLX-$\alpha$) operator (Ranjan et al., 2025), which generates offspring within an extended interval around the parent genes. Offspring $o=\{o_{1},o_{2},\ldots,o_{n}\}$ is generated from parents $x=\{x_{1},x_{2},\ldots,x_{n}\}$ and $y=\{y_{1},y_{2},\ldots,y_{n}\}$, where each $o_{i}$ is uniformly randomly chosen from the interval $\min(x_{i},y_{i})-\alpha D,\ \max(x_{i},y_{i})+\alpha D]$, with $D=|x_{i}-y_{i}|$ and $\alpha$ is constant.
@@ 2.4.3. Particle Swarm Optimization
We also implemented particle swarm optimization (PSO), another population-based meta-heuristic inspired by the social behavior of bird flocking or fish schooling (Bonyadi and Michalewicz, 2017). A swarm of $N$ particles is initialized, where each particle represents a complete UAV deployment solution $u=\{(x_{1},y_{1}),\allowbreak\ldots,\allowbreak(x_{u},y_{u})\}$. The initial position of each UAV for every particle is generated randomly within the search area $A_{r}$, identical to the initialization process described for the GA. Furthermore, each particle is assigned an initial velocity vector, typically set to zero or a small random value (Guo et al., 2025).
In each iteration, the velocity and position of each particle are updated based on its personal best known position ($p_{\text{best}}$) and the global best known position in the swarm ($g_{\text{best}}$) (Guo et al., 2025). The velocity $v_{i}$ and position $x_{i}$ of the $i$-th particle are updated as follows:
@@ 2.4.4. Simulated Annealing
Simulated Annealing (SA) is a global optimization algorithm inspired by the physical annealing process, a technique that involves heating and controlled cooling of a material to alter its physical properties (Kirkpatrick et al., 1983). SA operates by exploring new solutions in the neighborhood of the current solution (Khurshid and Maqsood, 2024). When a better solution is found, the algorithm moves toward it to improve performance. This process is repeated over multiple iterations, where at iteration $k$, the cooling parameter controls the acceptance of new solutions. In the early stages, a high cooling parameter allows the algorithm to accept a wider range of solutions and explore the search space broadly. As the iterations progress, the cooling parameter gradually decreases, enabling a more refined search and guiding the algorithm toward the final solution.
@@ 48    $d_{c}\leftarrow\text{HaversineDistance}(P_{repaired}[p],C_{center})$
We employed linear interpolation, which estimates intermediate values when the endpoints are known, to relocate UAVs that are outside the $A_{r}$ area back within the allowed region (Kholiavchenko et al., 2024; Scott et al., 2022). Specifically, given a UAV position $p_{repaired}$, the UAV control center $C_{\text{center}}$, and the allowed search area radius $R_{\text{search}}$, the repaired UAV position $p_{\text{repaired}}$ is computed as $p_{\text{repaired}}\leftarrow C_{\text{center}}+\frac{p_{repaired}-C_{\text{center}}}{\|p_{repaired}-C_{\text{center}}\|}\times R_{\text{search}}$. This formula computes the unit vector pointing from the center to the UAV and scales it by the search radius, ensuring that the UAV is placed exactly on the boundary. The same mechanism is used both for the initial correction and as a boundary clamping step after applying repulsive forces.
We then applied the virtual force algorithm (VFA) to minimize overlap between UAVs. In VFA, virtual forces such as attractive and repulsive forces are typically employed to control the movement of agents by drawing them toward targets and pushing them away from obstacles or other agents (Yoon and Kim, 2013). However, in our study, only the repulsive force $\vec{F}_{r}$ is used to prevent UAVs from overlapping. When the haversine distance between two UAVs, $u_{i}=(x_{i},y_{i})$ and $u_{j}=(x_{j},y_{j})$, is less than the sum of their detection radii, the positions of $u_{i}$ and $u_{j}$ are updated by adding $\alpha_{r}\cdot\frac{\vec{f}_{r}}{n_{r}}$, where $\vec{f}_{r}\leftarrow\vec{f}_{r}+\left(\frac{r_{u_{i}}+r_{u_{j}}}{d(u_{i},u_{j})}\right)\cdot(u_{j}-u_{i})$, $d(u_{i},u_{j})$ is the haversine distance between $u_{i}$ and $u_{j}$, $r_{u_{i}}$ and $r_{u_{j}}$ are their detection radii, $n_{r}$ is the number of repulsive interactions, and $\alpha_{r}$ is a constant weighting parameter.
@@ 3.2. Experimental setup
Table 1 lists the settings of parameters used in this study. The population size and the number of generations in the GA were both set to 50. The crossover and mutation rates were set to 100% and 10%, respectively. We used three instances to evaluate the proposed framework. For BLX-$\alpha$, the constant parameter $\alpha$ is set to 0.5, which has been widely used (Yoon and Kim, 2021; Tsutsui and Goldberg, 2001). In the repair phase of VFA, $\alpha_{r}$ is set to 0.9. Additionally, different settings—such as the number of particles and UAVs—were applied to assess the generalizability of the framework’s performance. During the GA process, a unit length of 100 (equivalent to 100 meters) was used to segment the straight lines between particles and the accident location. In other words, we employed a coarse approximation during the GA to optimize performance efficiently, followed by a detailed evaluation for accuracy After the GA phase, a finer unit length of 1 was used to compare the actual trajectories of the ocean drifters with the UAV deployment results, and the number of initial missing ocean drifters, $K_{0}$, was set to 100. For PSO, the inertia weight $w$ was set to 0.7, and both cognitive and social coefficients, $c_{1}$ and $c_{2}$, were set to 2.0. The SA algorithm was configured with an initial temperature of $T=1.0$, a cooling rate of $\alpha=0.95$, and a number of iterations was set to 10.
@@ 3.3.1. Prediction
To evaluate the predictive performance of the language model, we compared it with models previously employed for ocean drifter forecasting (Kim et al., 2025, 2024). To mimic real-world scenarios, we calculated the Haversine distance between the predicted and actual values. As shown in Table 2, while the conventional benchmark models outperformed the larger Chronos-base and Chronos-large variants, which had errors over 1.4 km, the compact Chronos-tiny model achieved superior performance, yielding a lower Haversine error (0.979 km) compared to the best-performing benchmark, Linear Regression (1.142 km).
@@ 6. Acknowledgments
This research was supported by Korea Institute of Marine Science & Technology Promotion (KIMST) funded by the Ministry of Oceans and Fisheries, Korea (RS-2022-KS221629). This research was also supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education(No. RS-2025-25419851)